OpenStudy (jane11509):

Identify the property shown. 6 • w = w • 6

OpenStudy (jane11509):

I just wanna work this problem out I don't want the straight answer

OpenStudy (igreen):

Properties of Multiplication: Commutative Property: \(a \cdot b = b \cdot a\) Associative Property: \((a \cdot b) \cdot c = a \cdot (b \cdot c)\) Identity Property: \(a \cdot 1 = a\)

OpenStudy (igreen):

Based on those properties of Multiplication, which do you think matches this one?

OpenStudy (jane11509):

Commutative Property Thank you ^_^